Sarah (single) purchased a home on January 1, 2008, for $600,000. She eventually sold the home for $800,000. What amount of the $200,000 gain on the sale does Sarah recognize in each of the following alternative situations? (Assume accumulated depreciation on the home is $0 at the time of the sale.) (Leave no answer blank. Enter zero if applicable.)

b. Sarah used the property as a vacation home through December 31, 2016. She then used the home as her principal residence from January 1, 2017, until she sold it on January 1, 2020. (Round intermediate percentage computation to 2 decimal places.)

c. Sarah used the home as a vacation home from January 1, 2008, until January 1, 2019. She used the home as her principal residence from January 1, 2019, until she sold it on January 1, 2020.

d. Sarah used the home as a vacation home from January 1, 2008, through December 31, 2013. She used the home as her principal residence from January 1, 2014, until she sold it on January 1, 2020. (Round intermediate percentage computation to 2 decimal places.)
Answer #1

Part B

Gain recognized

$133340

Home is sold after December 31, 2008 by Sarah and nonqualified use after December 31, 2008, she is not allowed to exclude a % of the gain that would otherwise be excluded.

The numerator in fraction of percentage of the gain that is not excluded is 8 (January 1, 2009 - December 31, 2016) and denominator is 12 (January 1, 2008 - January 1, 2020). Therefore, 8/12 = 66.67%

Gain recognized (gain that is not eligible for exclusion) = 200000*66.67% = $133340

Part C

Gain recognized

$200000

Sarah clears the ownership test, however she does not clear the use test as the use of property as principal residence was done for less than two of the five years preceding the sale

Part D

Gain recognized

$83340

Home is sold after December 31, 2008 by Sarah and nonqualified use after December 31, 2008, she is not allowed to exclude a % of the gain that would otherwise be excluded.

The numerator in fraction of percentage of the gain that is not excluded is 5 (January 1, 2009 - December 31, 2013) and denominator is 12 (January 1, 2008 - January 1, 2020). Therefore, 5/12 = 41.67%

Gain recognized (gain that is not eligible for exclusion) = 200000*41.67% = $83340